# Uncovering Grounding IDs: How External Cues Shape Multimodal Binding

### v3

@@ Abstract
Large vision–language models (LVLMs) show strong performance across multimodal benchmarks but remain limited in structured reasoning and precise grounding. Recent work has demonstrated that adding simple visual structures, such as partitions and annotations, improves accuracy, yet the internal mechanisms underlying these gains remain unclear. We investigate this phenomenon and propose the concept of Grounding IDs, latent identifiers induced by external cues that bind objects to their designated partitions across modalities. Through representation analysis, we find that these identifiers emerge as consistent within-partition alignment in embedding space and reduce the modality gap between image and text. Causal interventions further confirm that these identifiers mediate binding between objects and symbolic cues. We show that Grounding IDs strengthen attention between related components, which in turn improves cross-modal grounding and reduces hallucinations. Taken together, our results identify Grounding IDs as a key symbolic mechanism that explains how external cues enhance multimodal binding and offer both interpretability and practical improvements.
@@ 1 Introduction
Large vision–language models (LVLMs), such as LLaVA (llava), GPT-4V (gpt4), and Qwen-VL (qwen_vl) have achieved strong results in multimodal tasks, including image captioning, visual question answering, and embodied tasks. Despite this progress, they remain limited in structured reasoning (binding_problem), precise grounding, and hallucination (hallucination_survey). Very recent studies suggest that external visual structure can systematically improve LVLM reasoning. For instance, forgotten_polygons reveal that many LVLMs struggle with basic polygon identification, while adding shape annotations substantially boosts accuracy. Notably, VISER (viser) introduces more general visual structure (e.g., grids or horizontal lines) and demonstrates consistent gains across diverse reasoning tasks. Despite compelling advancements, the interpretability and mechanism-level understanding of this phenomenon remain underexplored. Prior work has largely focused on what visual structure can enhance reasoning, but not how it helps.
To bridge this critical understanding gap, we introduce Grounding IDs, symbolic multimodal identifiers induced by external cues. Our central hypothesis is that providing LVLMs with external visual cues (e.g., row labels, symbols, grid lines) aligned with textual cues (e.g., prompts for symbol-based scanning) leads LVLMs to generate latent identifiers that bind objects to their corresponding cues. These Grounding IDs propagate through embeddings and attention, enabling consistent correspondence within and across modalities. We validate this hypothesis through several analyses. First, observational probes show that Grounding IDs shape attention, reinforcing partition-level dynamics across modalities. Alignment experiments further reveal that symbolic cues strengthen within-partition binding and reduce the modality gap. Second, causal interventions confirm that external cues become causally bound to multimodal object tokens through Grounding IDs.
Beyond interpretability, we extend prior work on simple scaffolds (e.g., horizontal lines) to more effective multimodal cues, which align input modalities through a set of unique marks. Our ablation studies show that both visual and textual cues contribute to binding, with their combination yielding the greatest improvements in cross-modal grounding. Finally, we demonstrate the practical utility of Grounding IDs by showing that enhanced cross-modal attention substantially reduces hallucinations in LVLMs. Overall, our results identify Grounding IDs as a core mechanism of partition-based binding, providing both mechanistic insight into LVLM reasoning and also practical gains on visual reasoning tasks.
@@ 2 Background and Hypotheses
Earlier studies show that adding external artifacts, combined with chain-of-thought (CoT) style prompting, can substantially enhance the reasoning abilities of vision–language models by shifting them from one-pass perception toward more systematic, system-2-like processing. forgotten_polygons demonstrated that LVLMs are often “shape-blind”: their vision encoders cluster frequent shapes but fail to distinguish less common ones, leading to errors in side counting and geometric reasoning. Explicit cues, such as annotated edges, encourage more deliberate strategies and yield large accuracy gains.
VISER (viser) generalizes this idea beyond polygons by introducing versatile and input-agnostic visual structure, such as horizontal lines, paired with sequential scanning prompts that promote structured reasoning. This approach reduces feature binding errors, encourages serial scene parsing, and consistently improves performance across reasoning tasks like counting and visual search. Building on this foundation, our work probes the internal circuits through which simple external artifacts shape visual reasoning in LVLMs.
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that link entities with their attributes (binding_id). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they employ similar vectors linking visual objects with textual references (vision_binding_id), with properties such as factorizability and position independence. These works examine binding between items and attributes in the standard setting. Inspired by these abstract identifiers, we investigate the underlying mechanism by which external visual structures, equipped with aligned textual cues, enhance LVLMs’ reasoning.
@@
•  Existence: Augmenting inputs with multimodal external cues induces Grounding IDs that propagate through embeddings and attention, establishing within-partition binding across modalities.
@@
•  Causal Relation: Grounding IDs link objects to external cues in a causal manner, amplifying embedding alignment and increasing attention between connected parts.
To evaluate these hypotheses, the remainder of the paper is organized as follows. In Section 3, we probe model representations and internal dynamics to provide empirical evidence for Grounding IDs, showing how they shape attention patterns and reduce the modality gap. Section 4 presents causal intervention experiments that directly test whether Grounding IDs mediate object–cue binding. Finally, Section 5 explores the practical implications, demonstrating that enhancing cross-modal binding through Grounding IDs reduces hallucinations in LVLMs.
3 Empirical Evidence for Existence of Grounding IDs
Figure 2: Illustration of attention patterns under baseline and structured inputs in the scene description task. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices.
To investigate why external cues improve performance, we first analyze internal attention patterns of LVLMs. We focus on inference-time reasoning only, so we use a frozen 7B Qwen2.5-VL model without any fine-tuning. The tasks involve scene description and visual question answering, where the model produces detailed descriptions of the scene based on the provided input. The analysis is conducted on a synthetic dataset of 100 images. Each image contains 15 unique objects drawn from 35 shape–color combinations (7 shapes × 5 colors). To keep the analysis straightforward, each object occupies a single patch (28×28 pixels) and does not extend into adjacent patches. The baseline input follows the standard LVLM setup: the unmodified image, together with a plain scene-description prompt. The structured input augments this baseline by adding four symbols (&, #, \$, @) to the image and prompt, and by dividing the image into four horizontal partitions with three lines, as shown in Fig. 2(a). Ablations of alternative cue designs are provided in Appendix A.4.
For the representation analysis, we first determine the textual tokens that correspond to visual objects after the model generates its output. Text tokens are matched to visual objects using regular-expression (regex) pattern matching, since each object type appears only once in the image. This process assigns a partition label to each token in both modalities, allowing us to compute attention and embedding statistics in a partition-specific manner. Visual tokens are then grouped by their location within the four rows. Finally, we analyze attention patterns and embedding similarities within visual or cross-modal regions.
@@ 3 Empirical Evidence for Existence of Grounding IDs
For each token, we take the maximum attention score over all heads and then average across tokens within each partition (image rows), yielding a $4\times 4$ matrix. Fig. 2 shows both within-modality and cross-modality attention matrices for baseline and structured inputs. Structured inputs exhibit stronger diagonal dominance, with attention concentrated within partitions. The same trend appears in cross-modal attention between visual and textual tokens, where structured inputs achieve stronger cross-attention than baselines. These results indicate that external cues induce systematic partition-based binding that enhances the model’s attention.
@@
While attention captures binding structure, embedding similarity provides a complementary view of alignment between modalities. We measure cosine similarity between L2-normalized visual and textual embeddings corresponding to correctly generated object tokens. Similarities are computed layer-wise and averaged across samples. Both baseline and structured inputs show alignment strengthening in later layers (after layer 20). Structured inputs consistently achieve higher similarity, particularly in the last four layers, as shown in Fig. 3. This confirms that external cues reduce the modality gap in LVLMs by enhancing cross-modal alignment. In particular, cosine similarity between activation patches corresponding to external symbols across modalities is higher than that of dataset objects. modality_gap report that post-hoc removal of gap-contributing dimensions, while effective in reducing the modality gap in VLMs, distorts the embedding geometry and degrades performance. In contrast, our results show that external multimodal cues reduce the modality gap through the model’s internal dynamics without such disruptive interventions.
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, increasing alignment and grounding. We now seek causal validation using a medium-sized LVLM, Qwen2.5-VL 7B. For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols {&,\$,#,@} to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, “What is in row @?”, and the model must respond with the correct object description, such as “red circle.”
@@
•  $\mathbf{o}^{\sim s}_{s}$ (adjacent object with $s$): an object located in the partition of $s$ but bound to a different symbol.
•  $\mathbf{o}^{s}_{\sim s}$ (grounded object with $s$): an object bound to $s$ but located in a different partition.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to prior studies (vig2020investigating; binding_id). We adopt standard terminology from mechanistic interpretability to describe the causal experiments in this section (MI_Survey). We denote by $c$ an input context consisting of an image $x$ and a text prompt $p$. Given a model $M$ and layer $\ell$, let $h^{(\ell)}(c)$ be the hidden activations at that layer, including both visual patch tokens and textual tokens. In our interventions, we use three contexts: a target context $c$, a source context $c^{\prime}$, and a patched context $c^{*}$ obtained by replacing a subset of activations in $h^{(\ell)}(c)$ with those from $h^{(\ell)}(c^{\prime})$ and then running the remaining layers of $M$ on the modified activations.
@@
Two contexts, $c$ and $c^{\prime}$ are randomly sampled, and their activations are extracted across all layers. We then select two random symbols (e.g., \$ and #) and swap the patch activations of their corresponding object between the two contexts. For instance, the object in row \$ of $c^{\prime}$ is inserted into row # of $c$, while the object in row # of $c^{\prime}$ is patched into row \$. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are the same between $c$ and $c^{\prime}$. We intentionally choose the swap strategy to avoid ambiguity arising from the coexistence of $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{s}_{s}$.
@@
By analyzing the predictions of the patched context $c^{*}$, we discover that predictions consistently follow the binding symbols of the swapped objects rather than the symbols present locally in the host context. In other words, for a given partition $s$, the model outputs $\textbf{o}^{s}_{\sim s}$ (grounded object) instead of $\textbf{o}^{\sim s}_{s}$ (adjacent object). For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @ even though that object is physically located beside &. This indicates that the model disregards local cues and follows the transferred symbol–object binding for its final prediction (Fig. 4(a)).
@@
To quantify this effect across the dataset, we consider all valid swaps (across possible objects, symbols, and locations) and measure average swap accuracy. For each valid swap of object patches between $c$ and $c^{\prime}$, we evaluate the model on $c^{*}$ using a fixed set of query symbols $s\in\mathcal{S}$. If the model’s answer for symbol $s$ in $c^{*}$ matches the grounded object $\mathbf{o}^{s}_{\sim s}$ (i.e., it follows the transferred binding rather than the locally present object $\mathbf{o}^{\sim s}_{s}$), we count the trial as correct. Swap accuracy is the fraction of such queries for which the model outputs $\mathbf{o}^{s}_{\sim s}$.
@@
If we follow the standard setting without intervention and treat $\textbf{o}^{(\cdot)}_{s}$ as the correct label, standard accuracy drops sharply from $\mathbf{1.00}$ to $\mathbf{0.02}$ after swapping. In contrast, if we treat $\textbf{o}^{s}_{\sim s}$ as the correct label (i.e., prediction based on transferred bindings), swap accuracy remains high at $\mathbf{0.98}$, confirming that predictions follow Grounding IDs; thus, object–symbol bindings are causally mediated by these IDs. Fig. 4(b) shows the average log-probability of predicted objects before and after activation swapping.
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ across layers, showing where the model begins to favor the grounded object. (b) Signal-to-noise scores for attention differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ patches across heads and layers.

### v4

@@ Abstract
Large vision–language models (LVLMs) perform well on multimodal tasks, but their ability to reason and precisely align visual and textual information still has room for improvement. In this study, we show that external visual cues, such as symbols or grid lines, help LVLMs form more accurate connections between visual components, such as objects, and their corresponding textual descriptions, improving their grounding and reasoning abilities. We introduce the concept of Grounding IDs, which are latent identifiers that arise within the model as a result of external cues structuring both visual and textual modalities. Our analysis reveals that partition-inducing external cues lead to Grounding IDs that make better alignment between corresponding visual and text representations, helping the model focus on relevant information. We find that Grounding IDs enhance attention between related components, improving cross-modal grounding and reducing hallucinations. Overall, our results show that Grounding IDs are a key mechanism that enables external cues to improve cross-modal alignment, reduce errors, and enhance the overall performance of LVLMs across a range of multimodal tasks.
@@ 1 Introduction
Large vision–language models (LVLMs), such as LLaVA (Liu et al., 2023b), GPT-4V (Achiam et al., 2023), and Qwen-VL (Bai et al., 2023), have demonstrated strong performance on multimodal tasks like image captioning, visual question answering, and embodied tasks. However, these models still face significant challenges in aligning visual and textual information accurately, leading to hallucinations in generated text and limited visual reasoning capabilities. While recent studies, such as Rudman et al. (2025) and VISER (Izadi et al., 2025), have shown that adding external structures like shape annotations or grid lines can improve performance, the mechanisms behind these improvements and their effects on model performance remain unclear. This presents a critical gap in understanding how LVLMs leverage these cues to reduce errors and enhance task performance.
To address this gap, we investigate the internal mechanisms of LVLMs under externally induced structure in the visual and textual modalities, revealing the emergence of Grounding IDs. These are latent identifiers that the model generates to bind visual features, such as objects or spatial regions, to external cues like symbols or grid lines. Our central hypothesis is that when LVLMs are provided with such external structures, they create these Grounding IDs to improve the alignment between image and text. Through causal analyses, we demonstrate how these identifiers emerge within the model’s internal representations and how they propagate across embeddings. This finding clarifies how external structures improve the model’s ability to link visual and textual information, enhancing both grounding and reasoning. Our work provides new insights into the internal mechanisms of LVLMs, contributing to a deeper understanding of how these models process and connect multimodal information.
Furthermore, we extend prior work on simple scaffolds (e.g., horizontal lines) (Izadi et al., 2025) to more effective multimodal cues, which align input modalities through a set of unique marks incorporated in both visual and textual modalities. Our ablation studies show that both visual and textual cues contribute to binding, with their combination yielding the greatest improvements in cross-modal grounding. We demonstrate the practical utility of Grounding IDs by showing that enhanced cross-modal alignment reduces hallucinations in LVLMs and improves performance on visual reasoning tasks. Our results identify Grounding IDs as a key mechanism for partition-based binding, providing mechanistic insight into LVLM when visual and textual modalities are structured by aligned external cues.
@@ 2 Background and Hypotheses
Earlier studies show that adding external artifacts, combined with chain-of-thought (CoT) style prompting, can substantially enhance the reasoning abilities of vision–language models by shifting them from one-pass perception toward more systematic, system-2-like processing. Rudman et al. (2025) demonstrated that LVLMs are often “shape-blind”: their vision encoders cluster frequent shapes but fail to distinguish less common ones, leading to errors in side counting and geometric reasoning. Explicit cues, such as annotated edges, encourage more deliberate strategies and yield large accuracy gains.
VISER (Izadi et al., 2025) generalizes this idea beyond polygons by introducing versatile and input-agnostic visual structure, such as horizontal lines, paired with sequential scanning prompts that promote structured reasoning. This approach reduces feature binding errors, encourages serial scene parsing, and consistently improves performance across reasoning tasks like counting and visual search. Building on this foundation, our work probes the internal circuits through which simple external artifacts shape visual reasoning in LVLMs.
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that link entities with their attributes (Feng and Steinhardt, 2023). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they form similar vectors connecting visual objects with textual references (Saravanan et al., 2025). Notably, this study is limited to very simple images where grounding is trivial, and issues such as information loss or cross-modal misalignment do not arise. These works examine binding between items and attributes in the standard setting. However, we intend to investigate the underlying mechanism by which external visual structures, equipped with aligned textual cues, enhance LVLMs’ reasoning, particularly in more complex scenarios where grounding is non-trivial and cross-modal misalignment is more pronounced.
@@
  • Existence: Augmenting the inputs with multimodal external cues induces Grounding IDs that propagate through embeddings and attention, establishing within-partition binding across modalities.
@@
Figure 2: Illustration of attention patterns under baseline and structured inputs in the scene description task. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices. Values are averaged over 500 samples and layers 22–27 of Qwen2.5-VL.
To evaluate these hypotheses, the remainder of the paper is organized as follows. In Section 3, we probe model representations and internal dynamics to provide empirical evidence for Grounding IDs, showing how they emerge and influence attention patterns and cross-modal alignment. Section 4 presents causal intervention experiments that investigate how Grounding IDs contribute to object–cue binding. Finally, Section 5 explores the practical implications, demonstrating that enhancing cross-modal binding through Grounding IDs reduces hallucinations in LVLMs.
3 Evidence of Improved Alignment
In this section, we provide empirical evidence for the existence of Grounding IDs by analyzing the internal representations and attention patterns of LVLMs. Our goal is to investigate whether inducing structure by external cues, such as symbols and lines, leads to improving the alignment between visual and textual components. We focus on inference-time reasoning using a 7B Qwen2.5-VL model without fine-tuning. The tasks involve scene description and visual question answering, where the model generates detailed descriptions of the scene based on the provided input. Results for additional models are provided in Appendix N.
We compare two setups: a baseline and a structured input method. The baseline uses an unmodified image with a standard scene-description prompt. In contrast, the structured input method augments the baseline by adding four symbols (&, #, \$, @) to the image and prompt, and dividing the image into four horizontal partitions with three lines, as shown in Fig. 2(a). This structured input is designed to provide additional cues that guide the model in better aligning visual and textual information. To assess the impact of different cues, we also conduct ablation studies with alternative cue designs, which are detailed in Appendix E.
The analysis is conducted on a synthetic dataset of images with varying object configurations. Each image contains 15 unique objects drawn from 35 shape–color combinations (7 shapes × 5 colors), with each object occupying a single patch (28×28 pixels) and not extending into adjacent patches. After the model generates its output, we match textual tokens to visual objects using regular-expression (regex) pattern matching, as each object type appears only once in the image. This process assigns a partition label to each token in both modalities, allowing us to compute partition-wise attention and embedding statistics. We then analyze attention patterns and embedding similarities both within individual modalities and across the visual-textual interface. These analyses provide empirical evidence that structuring the image and generated text improves alignment between visual and textual components, suggesting the presence of a mechanism that more effectively links objects with their corresponding descriptions.
@@ 3 Evidence of Improved Alignment
For each token, we take the maximum attention score over all heads and then average across tokens within each partition (image rows), yielding a $4\times 4$ matrix. Aggregation is performed on true positive objects, where the model generates descriptions, and the corresponding objects are present in the image. Descriptions (including shape and color) are linked to the correct objects, ensuring accurate attention. To avoid ambiguity, the shapes in the image are unique, ensuring proper association between descriptions and objects. The aggregated attention score is averaged across true positives, providing a measure of cross-modal alignment between the model’s generated text and the objects in the image. Fig. 2 shows both within-modality and cross-modality attention matrices for baseline and structured inputs. Despite the simplicity of the external cues, structured inputs exhibit stronger diagonal dominance, with attention concentrated within partitions. See also Appendix B and C for complementary attention analysis.
@@
While attention captures binding structure, embedding similarity provides a complementary view of alignment between modalities. We measure cosine similarity between visual and textual embeddings corresponding to correctly generated object tokens. Similarities are computed layer-wise and averaged across samples. Both the baseline and the structured case show alignment strengthening in later layers (after layer 20). Structured modalities consistently achieve higher similarity, particularly in the last four layers, as shown in Fig. 3. This confirms that external cues reduce the modality gap in LVLMs by enhancing cross-modal alignment. In particular, cosine similarity between activation patches corresponding to external symbols across modalities is higher than that of dataset objects.
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, enhancing multimodal binding and thereby increasing alignment and grounding. We now seek causal validation using a medium-sized LVLM, Qwen2.5-VL 7B (see Appendix N for additional models). For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols {&,\$,#,@} to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, “What is in row @?”, and the model must respond with the correct object description, such as “red circle.”
@@
•  $\mathbf{o}^{\sim s}_{s}$: an object located in the partition of $s$ but bound to a different symbol than $s$.
•  $\mathbf{o}^{s}_{\sim s}$: an object bound to $s$ but located in a different partition than $s$.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to the introduced one in prior studies on language models (Vig et al., 2020; Feng and Steinhardt, 2023). We adopt standard terminology from mechanistic interpretability to describe the causal experiments in this section (Rai et al., 2024). We denote by $c$ an input context consisting of an image $x$ and a text prompt $p$. Given a model $M$ and layer $\ell$, let $h^{(\ell)}(c)$ be the hidden activations (residual stream) at that layer, including both visual patch tokens and textual tokens. In our interventions, we use three contexts: a target context $c$, a source context $c^{\prime}$, and a patched context $c^{*}$ obtained by replacing a subset of activations in $h^{(\ell)}(c)$ with those from $h^{(\ell)}(c^{\prime})$ and then running the remaining layers and tokens of $M$ on the patched context $c^{*}$.
@@
For our analysis, two contexts $c$ and $c^{\prime}$ are randomly sampled from the controlled dataset, and their activations are extracted across all layers (see Appendix L for details on the dataset). We then select two random symbols (e.g., & and @) and swap the patch activations corresponding to their objects between the two contexts. Specifically, the activations from all layers of the object in row & of $c^{\prime}$ are replaced by the activations of the corresponding patches in row @ of $c$, while the object activations in row @ of $c^{\prime}$ are patched into row &. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are originally assigned different symbols between $c$ and $c^{\prime}$. This swap strategy is designed to avoid ambiguity that could arise from objects being bound to the same symbol in the patched context (i.e., coexisting $\mathbf{o}^{s}_{s}$ and $\mathbf{o}^{s}_{\sim s}$).
@@
By analyzing the predictions of the patched context $c^{*}$, we observe that the model consistently follows the bounded symbols to the objects in the source rather than the symbols present locally in the host context. For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @, even though that object is physically located beside &. This indicates that the symbol within a partition are encoded in the embedding of objects belonging to that partition and are transferred to the patched context through the patched objects. Moreover, the model disregards local cues in the patched context $c^{*}$ and follows the transferred symbol-object binding of $c^{\prime}$ for its final prediction (Fig. 4(a)).
@@
To quantify this effect across the dataset, we consider all valid swaps, where object patches are swapped between two contexts, $c$ and $c^{\prime}$, ensuring that the objects differ in both shape and color between the contexts. We evaluate the model on the patched context, $c^{*}$, using a fixed set of query symbols $s\in\mathcal{S}$. If the model’s answer for symbol $s$ in $c^{*}$ matches $\mathbf{o}^{s}_{\sim s}$ (i.e., it follows the transferred binding rather than the locally present object $\mathbf{o}^{\sim s}_{s}$), we count the trial as correct. Swap accuracy is the fraction of such queries for which the model outputs $\mathbf{o}^{s}_{\sim s}$. For further details on how valid swaps are generated and the full procedure, please refer to Section L in the appendix.
@@
If we follow the standard setting without intervention and treat object located in the partition corresponding to symbol $s$ as the desired output, standard accuracy drops sharply from $\mathbf{1.00}$ to $\mathbf{0.02}$ after swapping. In contrast, if we treat the object bound to symbol s in the target context being transferred to the patched context as the desired output, swap accuracy remains high at $\mathbf{0.98}$, confirming that predictions follow Grounding IDs. Thus, object-symbol bindings are causally mediated by these IDs. Fig. 4(b) shows the average log-probability of predicted objects before and after activation swapping.
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\mathbf{o}^{s}_{\sim s}$ and $\mathbf{o}^{\sim s}_{s}$ across layers, showing where the model begins to favor the bound object $\mathbf{o}^{s}_{\sim s}$. (b) Signal-to-noise scores for attention differences between $\mathbf{o}^{s}_{\sim s}$ and $\mathbf{o}^{\sim s}_{s}$ patches across heads and layers.